A car is moving around a banked curve of radius 86 m at a constant speed. The road is banked at an angle of 11° and the car is moving at just the correct speed to ensure that there are no lateral frictional forces on the tyres - that is the only forces acting on the car are the force of gravity and the normal force from the road.





At what speed is the car moving? (in m s−1 to 2 s.f)

1
Expert's answer
2022-04-07T18:35:45-0400

"v=\sqrt{gr\tan\alpha}=12.8~\frac ms."